OpenStudy (anonymous):

Write the equation in standard form: 6 – 2(x – y) = 3x + 5

OpenStudy (anonymous):

A. 5x – 2y = 1 B. 5x – 2y = –1 C. 5x + 2y = 1 D. 2x – 5y = 1

OpenStudy (haseeb96):

6-2x+2y=3x+5 3x+2x-2y+5-6=0 5x-2y-1=0 5x-2y=1 A is corrrect